# Model-free High Dimensional Mediator Selection with False Discovery Rate Control

### v2

@@ 5 Discussions
Since we use data splitting, we have $\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent. Define p-values $p_{j}=1$ if $W_{j}<0$ and $p_{j}=1/2$ if $W_{j}>0$, then Lemma B1 guarantee that we have $p_{j}\geq\text{Uniform}[0,1]$ and are independent from nonnulls. Since we assumed that the path-a statistics satisfy the weak dependence assumption, we have $p_{j}$s are weakly correlated. Applying Lemma B3, we obtain the mFDR and FDR control results. ∎
@@
C.2: Additional Simulation Results
@@
Figure C1: Simulation results for the empirical power and false discovery rate (FDR) with sample size $n=1000$ are shown for correlated mediators with a defined correlation structure across 100 replications. The results are stratified by the effect size of the correlation structure ($\rho$), the number of potential mediators (p), the effect size of path-a (a), and the effect size of path-b (b) in the non-linear models with Setting 1 in the upper, Setting in the lower.
Figure C1 summarizes the empirical results for non-linear models with interactions (Setting 1) and models where mediators undergo a cosine transformation. (Setting 2). In Setting 1, we observe that GKMS consistently achieves the highest power across all examined conditions, particularly under low to moderate correlation levels ($\rho\leq 0.5$). As the correlation among mediators increases, the power of all methods declines; however, GKMS demonstrates a more gradual decrease and remains superior throughout. When varying the signal strengths $a$ (path-a) and $b$ (path-b), GKMS continues to dominate in terms of power, while maintaining FDR well below or near the nominal level of 0.1. In contrast, eHIMA often fails to control FDR, particularly under weak signals or high correlation, and shows substantially reduced power.

### v3

@@ 5 Discussions
Since we use data splitting, we have $\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent. Define p-values $p_{j}=1$ if $W_{j}<0$ and $p_{j}=1/2$ if $W_{j}>0$, then Lemma 1 guarantee that we have $p_{j}\geq\text{Uniform}[0,1]$ and are independent from nonnulls. Since we assumed that the path-a statistics satisfy the weak dependence assumption, we have $p_{j}$s are weakly correlated. Applying Lemma 3, we obtain the mFDR and FDR control results. ∎
@@
C.2: Details of Method specification
[…]
C.3: Additional Simulation Results
@@
Figure 2 summarizes the empirical results for non-linear models with interactions (Setting 1) and models where mediators undergo a cosine transformation. (Setting 2). In Setting 1, we observe that GKMS consistently achieves the highest power across all examined conditions, particularly under low to moderate correlation levels ($\rho\leq 0.5$). As the correlation among mediators increases, the power of all methods declines; however, GKMS demonstrates a more gradual decrease and remains superior throughout. When varying the signal strengths $a$ (path-a) and $b$ (path-b), GKMS continues to dominate in terms of power, while maintaining FDR well below or near the nominal level of 0.1. In contrast, eHIMA often fails to control FDR, particularly under weak signals or high correlation, and shows substantially reduced power.
@@
Appendix D.4: Additional Simulation Results
As a sensitivity analysis, we also applied our proposed method with Lasso linear regression in path-b (GKMS (Lasso)), HIMA, HIMA2, eHIMA, and MediationFDR methods for comparison. Results summarize in Tables 4, 5, and 6 for CDR-SB, ADAS-Cog, and MMSE scores, respectively. The mediation effects were estimated using structural equation modeling (SEM) through the lavaan R package (Rosseel, 2012).
Method Selected Mediator Effect (95% CI; p-value)
$\beta$ = -0.132 (-0.193; -0.071;<<0.001)
Mediation = 0.069 (0.034, 0.104; <<0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.178 (-0.238; -0.118;<<0.001)
Mediation = 0.09 (0.054, 0.126; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = -0.111 (-0.171; -0.052;<<0.001)
Mediation = 0.066 (0.029, 0.104; 0.001)
$\beta$ = 0.15 (0.089; 0.211;<<0.001)
Mediation = -0.072 (-0.106, -0.039; <<0.001)
Volume of Brain Stem $\alpha$ = -0.955(-1.061, -0.848; <<0.001)
$\beta$ = -0.121 (0.06; 0.183;<<0.001)
Mediation = -0.116 (-0.176, -0.056; <<0.001)
$\beta$ = -0.376 (-0.438; -0.314;<<0.001)
Mediation = 0.19 (0.139, 0.241; <<0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.282 (-0.343, -0.222; <<0.001)
Mediation = 0.143 (0.099, 0.186; <<0.001)
Volume of Inf.Lat.Vent (R) $\alpha$ = -0.483(-0.591, -0.376; <<0.001)
$\beta$ = 0.135 (0.074, 0.196; <<0.001)
Mediation = -0.065 (-0.098, -0.032; <<0.001)
ThickAvg2 of Parahip3 (R) $\alpha$ = 0.123 (0.007, 0.24; 0.037)
$\beta$ = -0.128 (-0.185, -0.072; <<0.001)
Mediation = -0.016 (-0.032, 0.001; 0.059)
$\beta$ = -0.376 (-0.438; -0.314;<<0.001)
Mediation = 0.19 (0.139, 0.241; <<0.001)
  • Abbreviations (all extracted from the aparc.a2009s statistics file):
  • 1 Inf.Lat.Vent: Inferior Lateral Ventricle.
  • 2 ThickAvg: Average cortical thickness of the specified cortical region.
  • 3 Parahip: Parahippocampal gyrus, a medial temporal lobe region.
Method Selected Mediator Effect (95% CI; p-value)
$\beta$ = -0.083 (-0.145, -0.02; 0.009)
Mediation = 0.043 (0.009, 0.077; 0.012)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.139 (-0.201, -0.078; <<0.001)
Mediation = 0.07 (0.036, 0.105; <<0.001)
ThickAvg1 of Angular Gyrus(L) $\alpha$ = 0.233 (0.118, 0.349; <<0.001)
$\beta$ = -0.099 (-0.157, -0.041; 0.001)
Mediation = -0.023 (-0.041, -0.005; 0.011)
ThickAvg of S.temporal.sup2(R) $\alpha$ = 0.092 (-0.023, 0.207; 0.117)
$\beta$ = -0.167 (-0.226, -0.109; <<0.001)
Mediation = -0.015 (-0.035, 0.005; 0.131)
$\beta$ = 0.081 (0.019, 0.143; 0.011)
Mediation = -0.058 (-0.104, -0.013; 0.012)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.324 (-0.387, -0.261; <<0.001)
Mediation = 0.164 (0.116, 0.211; <<0.001)
$\beta$ = -0.324 (-0.387, -0.261; <<0.001)
Mediation = 0.164 (0.116, 0.211; <<0.001)
$\beta$ = -0.063 (-0.123, -0.004; 0.036)
Mediation = 0.013 (-0.001, 0.027; 0.078)
ThickAvg of Pole temporal(R) $\alpha$ = 0.259 (0.143, 0.376; <<0.001)
$\beta$ = -0.084 (-0.146, -0.023; 0.007)
Mediation = -0.022 (-0.041, -0.003; 0.022)
  • We use version ADAS-13 representing the results of ADAS-Cog.
  • Abbreviations (all extracted from the aparc.a2009s statistics file):
  • 1 ThickAvg: Average cortical thickness of the specified cortical region.
  • 2 S.temporal.sup: Superior temporal sulcus.
Method Selected Mediator Effect (95% CI; p-value)
$\beta$ = 0.111 (0.048, 0.174; 0.001)
Mediation = -0.058 (-0.093, -0.023; 0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.184 (0.122, 0.246; <<0.001)
Mediation = -0.093 (-0.13, -0.056; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = 0.068 (0.007, 0.129; 0.028)
Mediation = -0.041 (-0.078, -0.004; 0.032)
$\beta$ = -0.088 (-0.15, -0.026; 0.006)
Mediation = 0.042 (0.011, 0.074; 0.008)
$\beta$ = -0.147 (-0.206, -0.088; <<0.001)
Mediation = 0.103 (0.059, 0.148; <<0.001)
$\beta$ = 0.336 (0.273, 0.399; <<0.001)
Mediation = -0.17 (-0.218, -0.121; <<0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.219 (0.158, 0.281; <<0.001)
Mediation = -0.111 (-0.15, -0.072; <<0.001)
ThickAvg2 of S.temporal.sup3 (L) $\alpha$ = 0.12 (0.007, 0.233; 0.037)
$\beta$ = 0.246 (0.187, 0.305; 0)
Mediation = 0.03 (0.001, 0.058; 0.043)
$\beta$ = 0.336 (0.273, 0.399; <<0.001)
Mediation = -0.17 (-0.218, -0.121; <<0.001)
  • Abbreviations (all extracted from the aparc.a2009s statistics file):
  • 1 Inf.Lat.Vent: Inferior Lateral Ventricle.
  • 2 ThickAvg: Average cortical thickness of the specified cortical region.
  • 3 S.temporal.sup: Sperior temporal sulcus.
Figure 2: Simulation results for the empirical power and false discovery rate (FDR) with sample size $n=1000$ are shown for correlated mediators with a defined correlation structure across 100 replications. The results are stratified by the effect size of the correlation structure ($\rho$), the number of potential mediators (p), the effect size of path-a (a), and the effect size of path-b (b) in the non-linear models with Setting 1 in the upper, Setting in the lower.